Find the derivative of the following functions (use the Rules of differentiation)


1.) y=(5x²-2x+1)²


2.) f(x)= cos (5x)


3. y=e exponent of 2x


4. g(x)= 10 exponent of x


1
Expert's answer
2021-12-01T16:34:10-0500

1.

"y'=2(5x^2-2x+1)(5(2x)-2)"

"=4(5x^2-2x+1)(5x-1)"

2.

"f'(x)= -5\sin (5x)"

3.

"y'=2e^{2x}"

4.


"g'(x)=\ln (10)\cdot10^x"